Give the equation of the line in standard form which is perpendicular to -3x+4y=8 and passes through (-3,5)
Answer by elima(1433)

3x+4y=8 and passes through (-3,5)
You need to first put the equation in slope-intercept form,
3x+4Y=8
4y=-3x+8
y=-3/4x+8/4
y=-3/4x+2
A line is perpendicular when its slopes are negative reciprocals of one another.
So the line which is perpendicular has a slope of 4/3.
So you need to plot the point given, then with the slope of 4/3 find the y-intercept then write the new equation.When you graph this line you find your y-intercept to be 1;
The equation would be;
y=4/3x+1
Hope you understand
=)